OpenStudy (anonymous):

Help reducing algebraic fractions? 12a^2-3/2 times (2a+1)^-2) times (6/(2a+1))^-1

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

12a^2-{3/2(1/(2a+1)^2*6(2a+1) =12a^2-3*3/(2a+1) =12a^2-9/(2a+1) =3(4a^2-3/(2a+1))